OpenStudy (anonymous):

Find the exact value of tan(arcsin(2/5)). For full credit, explain your reasoning.

OpenStudy (anonymous):

\[\arcsin(\frac{2}{5})=\theta\] is the same as \[\sin \theta=\frac{2}{5}\] so we draw a triangle sin is opposite over hypothenuse  solve for the missing side, then find \[\tan \theta\]

OpenStudy (anonymous):

soo 2/sqrt(21) is the answer?

OpenStudy (anonymous):

looks like it

OpenStudy (anonymous):

alright thanks